Two rainstorms occurred in one week in a certain area. In the first rainstorm 40ml of rain fell per hour, and in the second rainstorm 15ml of rain fell per hour. Rain fell that week for a total of 60 hours for a total rainfall of 1900ml . How long was each of the two rainstorms?

Answer by checkley71(8403)

15X+40(60-X)=1900 OR 15X+2400-40X=1900 OR -25X=-500 OR X=500/25 OR X=20 THUS IT RAINES 20 HOURS @ 15ML & (60-20)=40 HOURS @ 40ML
PROOF 15*20+40*40=1900 OR 300+1600=1900 OR 1900=1900